# Canonical decompositions of 3-connected graphs

### v2

@@ 1.7. Apex-decompositions
Assume Section 1.7. If $O$ has length at least four and $O$ does not have type btbt, then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations.
To prove Section 1.9 systematically, we need some machinery. Section 1.9 below will help us to find paths for verifying external tri-connectivity of mixed 3-separators. To make Section 1.9 applicable in various settings, we introduce the following definitions which will allow us to deal with cases systematically in Section 1.9 and its applications. Assume Section 1.7. A pattern is any of the following five words:
@@
Figure 1.10. The situation of Section 1.9 for $p={\tt btx}$ with ${\tt x}:={\tt b}$
@@
Assume Section 1.7. Suppose that $O$ has length at least four. Let $e_{1},e_{2},e_{3},e_{4}$ be consecutive edges in a cyclic orientation of $O$ with pattern $p$. Denote the endvertices of $e_{1}$ by $x_{0}$ and $x_{1}$ so that $x_{1}$ is incident with $e_{2}$.
@@
Proof of the Proof of Section 1.7: General case (1.9).
We consider each of the five possible patterns in turn. Let $G^{\prime}:=G-v$. Whenever an edge $e_{i}$ is bold, we let $B_{i}:=G^{\prime}[B_{\ell}]$ for the unique leaf-bag $B_{\ell}$ of ${\mathcal{A}}$ whose adhesion set consists of the endvertices of $e_{i}$ (tacitly assuming that $i$ is not a node of $S$). By Section 1.7, the vertex $v$ has a neighbour in $B_{i}\setminus O$ whenever $B_{i}$ exists, and we choose such a neighbour $v_{i}$ for each eligible $i$. We denote the vertex of $O$ that is incident with both $e_{i}$ and $e_{i+1}$ by $x_{i}$. This is consistent with the naming of $x_{1}$ in the statement of the lemma.
(bb) By Section 1.9, we find a 2-fan in $B_{1}$ from $x_{0}$ to $v_{1}$ and $x_{1}$. Since $B_{2}$ is 2-connected, we find an $x_{1}$$v_{2}$ path in $B_{2}$ which avoids $x_{2}$. The subgraph of $G$ obtained from the union of the 2-fan with the $x_{1}$$v_{2}$ path by adding $v$ and the edges $vv_{1}$ and $vv_{2}$ contains two desired paths.
(btx) By Section 1.9, we find a 2-fan in $B_{1}$ from $x_{0}$ to $v_{1}$ and $x_{1}$. If x is equal to b, then $B_{3}$ exists, and since $B_{3}$ is 2-connected, there is an $x_{2}$$v_{3}$ path in $B_{3}$ which avoids $x_{3}$. Then the subgraph of $G$ obtained from the union of the 2-fan with the path by adding the timid edge $e_{2}$ as well as $v$ and the two edges $vv_{1}$ and $vv_{3}$ contains two desired paths. Otherwise x is equal to t. Then the two edges $e_{2}$ and $e_{3}$ are timid, hence 3-connectivity implies that $x_{2}$ is adjacent to $v$. Then the subgraph of $G$ obtained from the 2-fan by adding the edges $vv_{1}$, $e_{2}$ and $vx_{2}$ contains two desired paths.
@@
Assume Section 1.7. Suppose that $O$ has length at least four. Let $\{a_{1},a_{2}\}$ be the adhesion set of a leaf-bag of ${\mathcal{A}}$ and let $i\in\{1,2\}$. If $a_{i}$ is not edgy, then either there are three internally disjoint paths from $a_{i}$ to $v$ avoiding $a_{3-i}$, or $va_{i}$ is an edge in $G$.
@@
Without loss of generality we have $i=2$. If $va_{2}$ is an edge in $G$ we are done, so let us suppose that $v$ is not adjacent to $a_{2}$. Let $e_{1},e_{2},e_{3}$ and $e_{4}$ be the four edges of $O$ that come after $a_{1}a_{2}$ on $O$ in the cyclic orientation of $O$ in which $a_{1}$ precedes $a_{2}$. Since $a_{2}$ is not edgy and $v$ is not a neighbour of $a_{2}$, the edge $e_{1}$ is bold. So the sequence $e_{1},e_{2},e_{3},e_{4}$ has pattern bb or btx, both of which have length at most three. Now we apply the Proof of Section 1.7: General case (1.9) to the sequence $e_{1},e_{2},e_{3},e_{4}$. This gives us two internally disjoint paths from $a_{2}$ to $v$ included in the reservoir. By assumption, $O$ has length at least four, so the vertex $a_{1}$ is distinct from the endvertices of the edges $e_{1}$ and $e_{2}$. Hence, the two internally disjoint paths avoid $B_{\ell}-a_{2}$, where $B_{\ell}$ is the leaf-bag with adhesion set $\{a_{1},a_{2}\}$. By Section 1.7, we find a third path from $a_{2}$ to $v$ included in $G[B_{\ell}+v]$, completing the proof. ∎
@@
Assume Section 1.7. Suppose that $O$ has length at least four and that $O$ does not have type btbt. Let $\{a_{1},a_{2}\}$ be the adhesion set of a leaf-bag of ${\mathcal{A}}$ and let $i\in\{1,2\}$. Denote the unique neighbour of $a_{i}$ on $O$ other than $a_{3-i}$ by $a_{i}^{\prime}$. If $a_{i}$ is edgy, then $a_{i}a_{i}^{\prime}$ is an edge in $G$ while $a_{i}v$ is not, and there are two internally disjoint paths from $a_{i}^{\prime}$ to $v$ in $G$ that avoid $B_{\ell}$.
@@
Without loss of generality, we have $i=2$. Let $e_{1},e_{2},e_{3}$ and $e_{4}$ be the four consecutive edges which come after $a_{1}a_{2}$ in the cyclic orientation of $O$ in which $a_{1}$ precedes $a_{2}$. Since $a_{2}$ is edgy and ${\mathcal{A}}$ is 2-connected, $e_{1}$ is timid. We apply the Proof of Section 1.7: General case (1.9) to the sequence $e_{1},e_{2},e_{3},e_{4}$. Since the pattern of this sequence starts with t and since $O$ has length at least four, we obtain two internally disjoint paths from $a_{2}^{\prime}$ to $v$ included in the reservoir and avoiding $a_{2}$.
If $O$ has length at least five, the vertex $a_{1}$ is distinct from the endvertices of the edges $e_{1},e_{2}$ and $e_{3}$. In particular, the two internally disjoint paths avoid the unique leaf-bag $B_{\ell}$ of ${\mathcal{A}}$ with adhesion set $\{a_{1},a_{2}\}$.
So it remains to consider the case that $O$ has length four. The existence of the bag $B_{\ell}$ implies that the edge $e_{4}$ is bold. In combination with our assumption that $O$ does not have the type btbt, it follows that the pattern tbtx is not possible (indeed, here ${\tt x}={\tt b}$ since $e_{4}$ is bold). Hence the pattern has length at most three. Thus the fact that the vertex $a_{1}$ is distinct from the endvertices of the edges $e_{1}$ and $e_{2}$ suffices to deduce that the two internally disjoint paths avoid $B_{\ell}$. This completes the proof. ∎
@@
Assume Section 1.7. If $O$ has length at least four and does not have type btbt, then $v$ has two neighbours in $G\setminus B_{\ell}$ for every leaf-bag $B_{\ell}$ of ${\mathcal{A}}$.
@@
Case $p={\tt bb}$. By Section 1.7, the vertex $v$ has two neighbours, one in $B_{1}\setminus O$ and one in $B_{2}\setminus O$.
Case $p={\tt btx}$. By Section 1.7, the vertex $v$ has one neighbour in $B_{1}\setminus O$. If the third edge is bold, we find a second neighbour in $B_{3}\setminus O$. Otherwise, we consider the endvertex that is shared by $e_{2}$ and $e_{3}$. By 3-connectivity, this endvertex must be adjacent to $v$. So $v$ has two neighbours outside of $B_{\ell}$.
@@
Assume Section 1.7. If $O$ has length at least four and does not have type btbt, then the tri-star of ${\mathcal{A}}$ consists of strong nontrivial tri-separations.
@@
Let $(X,Y)$ be a pseudo-reduction in the tri-star of ${\mathcal{A}}$, induced by a leaf $\ell$ of $S$. Let us denote the adhesion set of the leaf-bag $B_{\ell}$ by $\{a_{1},a_{2}\}$. We claim that very vertex $u\in S(X,Y)$ has degree at least four in $G$ and has at least two neighbours in both $X$ and $Y$.
@@
Case 2: $u=v$. Since $v$ lies in $S(X,Y)$, it has at least two neighbours in $B_{\ell}\subseteq X$. By Section 1.9, $v$ has two neighbours in $V(G)\setminus B_{\ell}\subseteq Y$.
Therefore, $(X,Y)$ is a strong tri-separation. It remains to show that $(X,Y)$ is nontrivial. Since $G^{\prime}[B_{\ell}]$ is 2-connected, it contains a cycle, which is included in $G[X]$. To see that $G[Y]$ contains a cycle, by Section 1.2 it suffices to show that $|Y\setminus X|\geqslant 2$, which follows from $O$ having length at least four. ∎
Proof of Section 1.9.
Assume Section 1.7. Further suppose that $O$ has length at least four and that $O$ does not have type btbt. We have to show that the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations. By Section 1.9, the tri-star of ${\mathcal{A}}$ consists of strong nontrivial tri-separations. So it remains to show that these are totally nested.
@@
By Claim 1.9.0.1 and Claim 1.9.0.2, $S(X,Y)$ is externally tri-connected. Since $(X,Y)$ also is half-connected, Section 1.5 gives that $(X,Y)$ is totally nested. ∎
Proof of the (Angry Tri-Separation Theorem). (1.1.1).
Let us assume for a contradiction that there exists a 3-connected graph $G$ that fails all three outcomes of Theorem 1.1.1, that is: all nontrivial tri-separations of $G$ are crossed; $G$ is neither a wheel nor a $K_{3,n}$ for any $n\geqslant 3$; and $G$ is not internally 4-connected. Then $G$ has a nontrivial strong tri-separation $(A,B)$ by Section 1.2, which is half-connected by Section 1.5. By assumption, the tri-separation $(A,B)$ is crossed by another tri-separation $(C,D)$. The tri-separation $(C,D)$ is nontrivial by Section 1.2. By Section 1.3 and using that $G$ is not a wheel such as $K_{4}$, all four links have size one, and the centre consists of a single vertex $v$. By Section 1.7, $G$ has a 2-connected apex-decomposition ${\mathcal{A}}$ with centre $v$, such that $(A,B)$ and $(C,D)$ interlace the tri-star of ${\mathcal{A}}$, and such that the central torso-cycle of ${\mathcal{A}}$ alternates between $S(A,B)-v$ and $S(C,D)-v$. As $G$ is not a wheel, ${\mathcal{A}}$ has at least one leaf-bag, and so the tri-star of ${\mathcal{A}}$ is non-empty. By Section 1.7, the tri-star of ${\mathcal{A}}$ consists of totally-nested nontrivial tri-separations, contradicting our assumption that all nontrivial tri-separations of $G$ are crossed. ∎
Chapter 2 Decomposing 3-connected graphs
2.1. Overview of this chapter
In this chapter, we prove the main result of the paper, Theorem 1. The proof of Theorem 1 offers additional structural insights, which lead us to a refinement of Theorem 1 that comes in the form of Theorem 2.2.1.
This chapter is organised as follows. In the next section we introduce the notation we need to then state Theorem 2.2.1. Like Theorem 1, this theorem will have three possible outcomes for the torsos, and we devote a section to the analysis of each possible outcome.
2.2. Basics
2.2.1. Generalised wheels
Figure 2.1. A concrete generalised wheel (left) and its apex-decomposition (right), where leaf-bags are indicated in blue and the centre plus its incident edges are red
The following definitions are supported by Figure 2.1. A $Y$-graph is a 3-star $K_{1,3}$ and the set of its 3 leaves is referred to as its attachment set. A concrete generalised wheel is a triple $(W,O,v)$ where $W$ is a graph obtained from a cycle $O$ and a vertex $v$ not on $O$ by doing the following, subject only to the condition that the resulting graph has minimum degree three:
@@ 2.2.1. Generalised wheels
We refer to $O$ as the rim of this concreted generalised wheel, and we refer to $v$ as its centre. For convenience, we write $W$ instead of $(W,O,v)$, and refer to $W$ as a concrete generalised wheel by a slight abuse of notation. Since concrete generalised wheels have minimum degree three, it is straightforward to show that they are 3-connected. The length of a concrete generalised wheel means the length of its rim.
@@
2.2.2. Splitting stars
Recall that a set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of (oriented) mixed-separations of $G$ is a star with leaves $A_{i}$ if $(A_{i},B_{i})<(B_{j},A_{j})$ for all distinct indices $i,j\in I$. We have seen in Section 1.7 that these stars naturally correspond to star-decompositions if they consist of separations only.
Let $S$ be a set mixed-separations of $G$. A star $\{\,(A_{i},B_{i}):i\in I\,\}\subseteq S$ with leaves $A_{i}$ is splitting if for every $(C,D)\in S$ there is $i\in I$ with either $(C,D)\leqslant(A_{i},B_{i})$ or $(D,C)\leqslant(A_{i},B_{i})$.

### v3

@@ 10 Proof of Section 8: General case
Assume Section 8. If $O$ has length at least four and $O$ does not have type btbt, then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations.
To prove Section 10 systematically, we need some machinery. Section 10 below will help us to find paths for verifying external tri-connectivity of mixed 3-separators. To make Section 10 applicable in various settings, we introduce the following definitions which will allow us to deal with cases systematically in Section 10 and its applications. Assume Section 8. A pattern is any of the following five words:
@@
Figure 13: The situation of Section 10 for $p={\tt btx}$ with ${\tt x}:={\tt b}$
@@
Assume Section 8. Suppose that $O$ has length at least four. Let $e_{1},e_{2},e_{3},e_{4}$ be consecutive edges in a cyclic orientation of $O$ with pattern $p$. Denote the endvertices of $e_{1}$ by $x_{0}$ and $x_{1}$ so that $x_{1}$ is incident with $e_{2}$.
@@
Proof of the Proof of Section 8: General case (10).
We consider each of the five possible patterns in turn. Let $G^{\prime}:=G-v$. Whenever an edge $e_{i}$ is bold, we let $B_{i}:=G^{\prime}[B_{\ell}]$ for the unique leaf-bag $B_{\ell}$ of ${\mathcal{A}}$ whose adhesion set consists of the endvertices of $e_{i}$ (tacitly assuming that $i$ is not a node of $S$). By Section 8, the vertex $v$ has a neighbour in $B_{i}\setminus O$ whenever $B_{i}$ exists, and we choose such a neighbour $v_{i}$ for each eligible $i$. We denote the vertex of $O$ that is incident with both $e_{i}$ and $e_{i+1}$ by $x_{i}$. This is consistent with the naming of $x_{1}$ in the statement of the lemma.
(bb) By Section 10, we find a 2-fan in $B_{1}$ from $x_{0}$ to $v_{1}$ and $x_{1}$. Since $B_{2}$ is 2-connected, we find an $x_{1}$$v_{2}$ path in $B_{2}$ which avoids $x_{2}$. The subgraph of $G$ obtained from the union of the 2-fan with the $x_{1}$$v_{2}$ path by adding $v$ and the edges $vv_{1}$ and $vv_{2}$ contains two desired paths.
(btx) By Section 10, we find a 2-fan in $B_{1}$ from $x_{0}$ to $v_{1}$ and $x_{1}$. If x is equal to b, then $B_{3}$ exists, and since $B_{3}$ is 2-connected, there is an $x_{2}$$v_{3}$ path in $B_{3}$ which avoids $x_{3}$. Then the subgraph of $G$ obtained from the union of the 2-fan with the path by adding the timid edge $e_{2}$ as well as $v$ and the two edges $vv_{1}$ and $vv_{3}$ contains two desired paths. Otherwise x is equal to t. Then the two edges $e_{2}$ and $e_{3}$ are timid, hence 3-connectivity implies that $x_{2}$ is adjacent to $v$. Then the subgraph of $G$ obtained from the 2-fan by adding the edges $vv_{1}$, $e_{2}$ and $vx_{2}$ contains two desired paths.
@@
Assume Section 8. Suppose that $O$ has length at least four. Let $\{a_{1},a_{2}\}$ be the adhesion set of a leaf-bag of ${\mathcal{A}}$ and let $i\in\{1,2\}$. If $a_{i}$ is not edgy, then either there are three internally disjoint paths from $a_{i}$ to $v$ avoiding $a_{3-i}$, or $va_{i}$ is an edge in $G$.
@@
Without loss of generality we have $i=2$. If $va_{2}$ is an edge in $G$ we are done, so let us suppose that $v$ is not adjacent to $a_{2}$. Let $e_{1},e_{2},e_{3}$ and $e_{4}$ be the four edges of $O$ that come after $a_{1}a_{2}$ on $O$ in the cyclic orientation of $O$ in which $a_{1}$ precedes $a_{2}$. Since $a_{2}$ is not edgy and $v$ is not a neighbour of $a_{2}$, the edge $e_{1}$ is bold. So the sequence $e_{1},e_{2},e_{3},e_{4}$ has pattern bb or btx, both of which have length at most three. Now we apply the Proof of Section 8: General case (10) to the sequence $e_{1},e_{2},e_{3},e_{4}$. This gives us two internally disjoint paths from $a_{2}$ to $v$ included in the reservoir of $e_{1},\ldots,e_{3}$. By assumption, $O$ has length at least four, so the vertex $a_{1}$ is distinct from the endvertices of the edges $e_{1}$ and $e_{2}$. Hence, the two internally disjoint paths avoid $B_{\ell}-a_{2}$, where $B_{\ell}$ is the leaf-bag with adhesion set $\{a_{1},a_{2}\}$. By Section 8, we find a third path from $a_{2}$ to $v$ included in $G[B_{\ell}+v]$, completing the proof. ∎
@@
Assume Section 8. Suppose that $O$ has length at least four and that $O$ does not have type btbt. Let $\{a_{1},a_{2}\}$ be the adhesion set of a leaf-bag of ${\mathcal{A}}$ and let $i\in\{1,2\}$. Denote the unique neighbour of $a_{i}$ on $O$ other than $a_{3-i}$ by $a_{i}^{\prime}$. If $a_{i}$ is edgy, then $a_{i}a_{i}^{\prime}$ is an edge in $G$ while $a_{i}v$ is not, and there are two internally disjoint paths from $a_{i}^{\prime}$ to $v$ in $G$ that avoid $B_{\ell}$.
@@
Without loss of generality, we have $i=2$. Let $e_{1},e_{2},e_{3}$ and $e_{4}$ be the four consecutive edges which come after $a_{1}a_{2}$ in the cyclic orientation of $O$ in which $a_{1}$ precedes $a_{2}$. Since $a_{2}$ is edgy and ${\mathcal{A}}$ is 2-connected, $e_{1}$ is timid. We apply the Proof of Section 8: General case (10) to the sequence $e_{1},e_{2},e_{3},e_{4}$. Since the pattern of this sequence starts with t and since $O$ has length at least four, we obtain two internally disjoint paths $P,Q$ from $a_{2}^{\prime}$ to $v$ included in the reservoir of $e_{2},\ldots,e_{4}$ and avoiding $a_{2}$.
If $O$ has length at least five, the vertex $a_{1}$ is distinct from the endvertices of the edges $e_{1},e_{2}$ and $e_{3}$. In particular, the two paths $P,Q$ avoid the unique leaf-bag $B_{\ell}$ of ${\mathcal{A}}$ with adhesion set $\{a_{1},a_{2}\}$.
So it remains to consider the case that $O$ has length four. The existence of the bag $B_{\ell}$ implies that the edge $e_{4}$ is bold. In combination with our assumption that $O$ does not have the type btbt, it follows that the pattern tbtx is not possible (indeed, here ${\tt x}={\tt b}$ since $e_{4}$ is bold). Hence the pattern has length at most three. So $P$ and $Q$ are contained in the reservoir of $e_{2},e_{3}$ and avoid $a_{1},a_{2}$, by the earlier application of the Proof of Section 8: General case (10) and since the 3-letter pattern starts with t for the edge $e_{1}$. In particular, $P$ and $Q$ avoid $B_{\ell}$. This completes the proof. ∎
@@
Assume Section 8. If $O$ has length at least four and does not have type btbt, then $v$ has two neighbours in $G\setminus B_{\ell}$ for every leaf-bag $B_{\ell}$ of ${\mathcal{A}}$.
@@
Case $p={\tt bb}$. By Section 8, the vertex $v$ has two neighbours, one in $B_{1}\setminus O$ and one in $B_{2}\setminus O$.
Case $p={\tt btx}$. By Section 8, the vertex $v$ has one neighbour in $B_{1}\setminus O$. If the third edge is bold, we find a second neighbour in $B_{3}\setminus O$. Otherwise, we consider the endvertex that is shared by $e_{2}$ and $e_{3}$. By 3-connectivity, this endvertex must be adjacent to $v$. So $v$ has two neighbours outside of $B_{\ell}$.
@@
Assume Section 8. If $O$ has length at least four and does not have type btbt, then the tri-star of ${\mathcal{A}}$ consists of strong nontrivial tri-separations.
@@
Let $(X,Y)$ be an arbitrary pseudo-reduction in the tri-star of ${\mathcal{A}}$, induced by a leaf $\ell$ of $S$ (if no such $(X,Y)$ exists, then the tri-star is empty and we are done). Let us denote the adhesion set of the leaf-bag $B_{\ell}$ by $\{a_{1},a_{2}\}$. We claim that very vertex $u\in S(X,Y)$ has degree at least four in $G$ and has at least two neighbours in both $X$ and $Y$.
@@
Case 2: $u=v$. Since $v$ lies in $S(X,Y)$, it has at least two neighbours in $B_{\ell}\subseteq X$. By Section 10, $v$ has two neighbours in $V(G)\setminus B_{\ell}\subseteq Y$.
Therefore, $(X,Y)$ is a strong tri-separation. It remains to show that $(X,Y)$ is nontrivial. Since $G^{\prime}[B_{\ell}]$ is 2-connected, it contains a cycle, which is included in $G[X]$. To see that $G[Y]$ contains a cycle, by Section 3 it suffices to show that $|Y\setminus X|\geqslant 2$, which follows from $O$ having length at least four. ∎
Proof of Section 10.
Assume Section 8. Further suppose that $O$ has length at least four and that $O$ does not have type btbt. We have to show that the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations. By Section 10, the tri-star of ${\mathcal{A}}$ consists of strong nontrivial tri-separations. So it remains to show that these are totally nested.
@@
By Claim 10.0.1 and Claim 10.0.2, $S(X,Y)$ is externally tri-connected. Since $(X,Y)$ also is half-connected, Section 6 gives that $(X,Y)$ is totally nested. ∎
Proof of Section 8.
Proof of the (Angry Tri-Separation Theorem). (1.1).
Let us assume for a contradiction that there exists a 3-connected graph $G$ that fails all three outcomes of Theorem 1.1, that is: all nontrivial tri-separations of $G$ are crossed; $G$ is neither a wheel nor a $K_{3,n}$ for any $n\geqslant 3$; and $G$ is not internally 4-connected. Then $G$ has a nontrivial strong tri-separation $(A,B)$ by Section 3, which is half-connected by Section 6. By assumption, the tri-separation $(A,B)$ is crossed by another tri-separation $(C,D)$. The tri-separation $(C,D)$ is nontrivial by Section 3. By Section 4 and using that $G$ is not a wheel such as $K_{4}$, all four links have size one, and the centre consists of a single vertex $v$. By Section 8, $G$ has a 2-connected apex-decomposition ${\mathcal{A}}$ with centre $v$, such that $(A,B)$ and $(C,D)$ interlace the tri-star of ${\mathcal{A}}$, and such that the central torso-cycle of ${\mathcal{A}}$ alternates between $S(A,B)-v$ and $S(C,D)-v$. As $G$ is not a wheel, ${\mathcal{A}}$ has at least one leaf-bag, and so the tri-star of ${\mathcal{A}}$ is non-empty. By Section 8, the tri-star of ${\mathcal{A}}$ consists of totally-nested nontrivial tri-separations, contradicting our assumption that all nontrivial tri-separations of $G$ are crossed. ∎
Chapter 2
Decomposing 3-connected graphs
1 Overview of this chapter
In this chapter, we prove the main result of the paper, Theorem 1. The proof of Theorem 1 offers additional structural insights, which lead us to a refinement of Theorem 1 that comes in the form of Theorem 2.1.
This chapter is organised as follows. In the next section we introduce the notation we need to then state Theorem 2.1. Like Theorem 1, this theorem will have three possible outcomes for the torsos, and we devote a section to the analysis of each possible outcome.
2 Basics
1 Generalised wheels
Figure 14: A concrete generalised wheel (left) and its apex-decomposition (right), where leaf-bags are indicated in blue and the centre plus its incident edges are red
The following definitions are supported by Figure 14. A $Y$-graph is a 3-star $K_{1,3}$ and the set of its 3 leaves is referred to as its attachment set. A concrete generalised wheel is a triple $(W,O,v)$ where $W$ is a graph obtained from a cycle $O$ and a vertex $v$ not on $O$ by doing the following, subject only to the condition that the resulting graph has minimum degree three:
@@ 1 Generalised wheels
We refer to $O$ as the rim of this concrete generalised wheel, and we refer to $v$ as its centre. For convenience, we write $W$ instead of $(W,O,v)$, and refer to $W$ as a concrete generalised wheel by a slight abuse of notation. Since concrete generalised wheels have minimum degree three, it is straightforward to show that they are 3-connected. The length of a concrete generalised wheel means the length of its rim.
@@
2 Splitting stars
Recall that a set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of (oriented) mixed-separations of $G$ is a star with leaves $A_{i}$ if $(A_{i},B_{i})<(B_{j},A_{j})$ for all distinct indices $i,j\in I$. We have seen in Section 8 that these stars naturally correspond to star-decompositions if they consist of separations only.
Let $S$ be a set of mixed-separations of $G$. A star $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}\subseteq S$ with leaves $A_{i}$ is splitting if for every $(C,D)\in S$ there is $i\in I$ with either $(C,D)\leqslant(A_{i},B_{i})$ or $(D,C)\leqslant(A_{i},B_{i})$.